### v1

# On Softmax Direct Preference Optimization for Recommendation

Abstract
Recommender systems aim to predict personalized rankings based on user preference data. With the rise of Language Models (LMs), LM-based recommenders have been widely explored due to their extensive world knowledge and powerful reasoning abilities. Most of the LM-based recommenders convert historical interactions into language prompts, pairing with a positive item as the target response and fine-tuning LM with a language modeling loss. However, the current objective fails to fully leverage preference data and is not optimized for personalized ranking tasks, which hinders the performance of LM-based recommenders. Inspired by the current advancement of Direct Preference Optimization (DPO) in human preference alignment and the success of softmax loss in recommendations, we propose Softmax-DPO (S-DPO) to instill ranking information into the LM to help LM-based recommenders distinguish preferred items from negatives, rather than solely focusing on positives. Specifically, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for LM-based recommenders, connected to softmax sampling strategies. Theoretically, we bridge S-DPO with the softmax loss over negative sampling and find that it has a side effect of mining hard negatives, which assures its exceptional capabilities in recommendation tasks. Empirically, extensive experiments conducted on three real-world datasets demonstrate the superiority of S-DPO to effectively model user preference and further boost recommendation performance while mitigating the data likelihood decline issue of DPO. Our codes are available at https://github.com/chenyuxin1999/S-DPO.
1 Introduction
Recommender systems aim to predict personalized rankings based on user preference data, i.e., historical interactions such as purchases, clicks, and ratings [1, 2]. Recently, leveraging the extensive world knowledge and powerful reasoning abilities of language models (LMs) [3, 4, 5], LM-based recommenders have been broadly explored [6, 7, 8]. These recommenders convert historical interaction data into language prompts and either perform in-context learning or fine-tune LMs, demonstrating notable advantages, including zero-shot and few-shot reasoning [9, 10, 11, 12], enhanced generalization abilities [13, 14], and rich semantic understanding [15, 16, 17, 18]. However, current LM-based recommenders typically utilize language modeling loss for personalized ranking objectives—predicting the next token—which significantly differs from the objective of modeling user preferences in recommendation tasks [19, 20].
We argue that the current objective of LM-based recommenders does not fully utilize preference data and is not optimized for personalized ranking tasks, thereby hindering recommendation performance. Most LM-based recommenders address recommendation tasks by leveraging specialized language prompts [13, 16, 21, 22, 23], incorporating collaborative signals as a new modality [17, 18, 24], or extending the vocabulary of LMs with item tokens [15, 25, 26, 27, 28, 29]. Typically, these recommenders pair each language prompt, including the user’s historical interaction item lists, with a single positive item and then update LM parameters using language modeling loss [13, 18]. Despite being designed for recommendation tasks, these LM-based recommenders do not consider negative items and are not directly optimized for personalized rankings. Such a training paradigm fails to fully leverage user preference data and overlooks the role of negative items in recommendations, thereby impeding the alignment of LMs with user preferences.
Inspired by the success of using human-labeled data to align LMs with human preferences [30, 31, 32] and advancements in direct preference optimization (DPO) [33, 34, 35], we make progress on aligning LMs with recommendations by fine-tuning them to predict the next item in accordance with the user’s preference. This preference alignment stage aims to instill ranking information into the LMs and help recommenders distinguish preferred items from negatives, rather than solely focusing on positives.
Towards this end, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for recommendation, connected to softmax sampling strategies [19, 36, 37, 38], which we call S-DPO. Specifically, we first devise supervised fine-tuning to inject domain knowledge and improve LM’s ability to follow the instructions before preference alignment phase, following [13, 32]. In the preference alignment stage, instead of constructing solely positive pairs, we initially pair each language prompt with both positive and randomly sampled multiple negatives to build text-based preference data. Building upon these preference data, we extend conventional DPO with the Bradley-Terry preference model [33, 39] on pairwise data to the Plackett-Luce preference model [40, 41], which handles relative rankings in recommendation tasks.
Benefiting from the use of multiple negatives in preference data, our S-DPO offers three appealing properties. On the one hand, S-DPO serves as the first specialized personalized ranking loss for LM-based recommenders, effectively utilizing multiple negatives and acknowledging the importance of preference data. Empirically, we demonstrate that it provides more effective ranking gradients and mitigates the instability associated with DPO training (cf. Section 4.2). On the other hand, we theoretically bridge the DPO loss with the traditional BPR loss [42] over pairwise data and connect S-DPO with the softmax loss over negative sampling (also known as contrastive loss in self-supervised recommendations, which achieves state-of-the-art performance [36, 43, 44]). This connection naturally underscores the ranking performance of S-DPO and highlights the critical role of multiple negatives. Furthermore, we find that S-DPO has a side effect of mining hard negative examples similar to contrastive learning paradigm [37], which not only boosts the performance but also accelerates the training process (cf. Section 3.1), assuring its exceptional capabilities in recommendation tasks.
Overall, our contributions can be concluded as follows:
  • We are among the first to point out that the widely used language modeling loss in LM-based recommendation is not designed for ranking tasks and fails to fully utilize user preference data, thereby hindering recommendation performance.
  • We propose S-DPO, an alternative version of DPO loss tailored for LM-based recommenders, incorporating multiple negatives to instill ranking information into LM.
  • We theoretically bridge S-DPO with the softmax loss over negative sampling to highlight the critical role of multiple negatives and find its side effect of mining hard negatives, assuring its capabilities.
Figure 1: Framework of S-DPO. Different from existing methods which fine-tune LMs with a language modeling loss without tailoring for recommendations, S-DPO proposes to explicitly instill ranking information into LMs. To take one step further, S-DPO incorporates multiple negatives in user preference data and generalizes pairwise DPO loss to softmax ranking loss.
2 Preliminary
In this section, we first formalize sequential recommendation as the task of aligning language models (LMs) with user preferences. Then, we discuss the general framework of current LM-based recommenders that utilizes language modeling loss to fine-tune LMs. Finally, we outline the training process widely used to align LMs with human preferences, including reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO).
Task Formulation.
Given the historical interactions $\mathcal{H}_{u}$ of one user $u$ in chronological order, the goal of LM-based sequential recommender $\mathcal{M}_{\theta}$, where $\theta$ represents trainable parameters, is to select the item $i_{p}$ preferred by user $u$ from candidate set $C=\{i_{j}\}_{j=1}^{N}$, where N is the number of candidates. This task requires that item $i_{p}$ be preferred over the other candidate items, denoted by $\mathcal{I}_{d}=C\backslash\{i_{p}\}$. This requirement explicitly defines a multi-negative preference understanding for LM-based recommenders, which can be formulated as follows:
$\forall i_{d}\in\mathcal{I}_{d},\quad i_{p}>_{u}i_{d},$ (1)
wherein $>_{u}$ stands for the preference of user $u$.
Fine-tuning LM-based recommenders.
Current LM-based recommenders widely adopt supervised fine-tuning (SFT) [32] on recommendation-specific data to enhance their performance [6, 8]. Generally, this involves two steps: structuring recommendation data as text-based pairs and then fine-tuning LMs based on these pairs. In the first step, for user $u$, a recommendation task prompt $x_{u}$ encompasses the user’s historical interactions $\mathcal{H}_{u}$, the candidate item set $C$, and a description of the sequential recommendation task. This prompt $x_{u}$ is paired with the title of the preferred item $i_{p}$ in the candidate set $C$, denoted as $e_{p}$, to form the pair data $(x_{u},e_{p})$. In the second step, the $(x_{u},e_{p})$ pairs are utilized to fine-tune the LM-based recommender $\mathcal{M}_{\theta}$ through language modeling loss. This loss, commonly used in SFT in language modeling tasks, implicitly treats the recommendation task as predicting the next token based on preceding tokens. Formally, the objective of optimizing the LM-based recommender $\mathcal{M}_{\theta}$ with pair data $(x_{u},e_{p})$ can be formulated as:
[…]
where $|e_{p}|$ is the number of tokens in $e_{p}$, $(e_{p})_{t}$ is the $t$-th token of $e_{p}$ and $(e_{p})_{<t}$ is the tokens preceding $(e_{p})_{t}$.
However, recommendation tasks are essentially user preference alignment tasks, as formalized in the above task formulation, and differ from user modeling tasks that consider only positive responses. Such a gap necessitates further exploration into aligning LM-based recommenders with user preference, an area that has been underexplored.
RLHF pipeline and DPO.
Recent studies in natural language processing (NLP) explore the use of human-labeled pairwise data as a reward signal to align LMs with human preferences, such as RLHF [32] and DPO [33]. Specifically, the RLHF [32] pipeline adds two additional phases after the SFT phase: reward model training and reinforcement learning (RL) optimization. After obtaining the SFT model $\pi^{\rm SFT}$, RLHF further optimizes it with pairwise preference data.
Inspired by the success of RLHF in NLP, we leverage RLHF to inject recommendation-specific user pairwise preference into LM-based recommenders. Let =$\mathcal{E}=\{e_{j}\}_{j=1}^{N}$ denote the title set of candidate items, where $e_{j}$ denotes the title of item $i_{j}$. Given two items $i_{j},i_{k}\in\mathcal{C}$, the user preference $i_{j}>_{u}i_{k}$ can be seamlessly transformed into a response preference, stipulating that $e_{j}$ is preferred over $e_{k}$ when presented with prompt $x_{u}$, denoted as $e_{j}\succ e_{k}|x_{u}$. By sampling one dispreferred item $i_{d}$ from dispreferred candidate set $\mathcal{I}_{d}$, we can curate a preference dataset $\{(e_{p},e_{d},x_{u})\}$.
After that, RLHF utilizes a preference model for preference distribution modeling, such as Bradley-Terry (BT) model [39]. This preference model assumes there is a latent function $r(x_{u},e_{j})$ representing the reward of prompt-response pair $(x_{u},e_{j})$. The bigger reward $r(x_{u},e_{j})$ means the more user $u$ prefers item $i$. From this perspective, reward function $r(x_{u},e_{j})$ serves as a scoring function that quantifies the preference of user $u$ to item $i$. Besides, the preference model defines a mapping from the reward function $r(x_{u},e_{j})$ to a preference distribution $p_{r}(e_{j}\succ e_{k}|x_{u})$. Based on preference distribution, an optimal reward function is trained by maximizing the likelihood of preference data. The training objective of this phase is as follows:
$\mathcal{L}_{\rm RM}=-\mathbb{E}_{(x_{u},e_{p},e_{d})}[{\rm log}\,p_{r}(e_{p}\succ e_{d}|x_{u})].$ (3)
Let $\pi_{\theta}(e|x_{u})$ be the probability that LM-based recommender $\mathcal{M}_{\theta}$ output title $e$ given prompt $x_{u}$. The final reinforcement learning phase aims to maximize the expected reward of policy while not deviate too far from the reference model, formulating the following objective for optimal policy:
$\underset{\pi_{\theta}}{\rm max}\,\mathbb{E}_{x_{u}\sim\mathcal{D},e\sim\pi_{\theta}(e|x_{u})}[r(x_{u},e)]-\beta\mathbb{D}_{\rm KL}[\pi_{\theta}(e|x_{u})||\pi_{\rm ref}(e|x_{u})],$ (4)
where $\mathcal{D}$ denotes the distribution of $x_{u}$ and $\pi_{\rm ref}=\pi^{\rm SFT}$.
A recent study, DPO [33], theoretically proves the optimal policy in a closed form to Eq.(4) is
$\pi^{*}(e|x_{u})=\frac{1}{Z(x_{u})}\pi_{\rm ref}(e|x_{u}){\rm exp}\left(\frac{1}{\beta}r(x_{u},e)\right),$ (5)
which is equivalent to
$r(x_{u},e)=\beta\,{\rm log}\,\frac{\pi(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}+\beta\,{\rm log}\,Z(x_{u}),$ (6)
where $Z(x_{u})=\sum_{e}\pi_{\rm ref}(e|x_{u}){\rm exp}\left(\frac{1}{\beta}r(x_{u},e)\right)$ is the partition function.
By defining $p_{r}(e_{p}\succ e_{d}|x_{u})$ as $\sigma(r(x_{u},e_{p})-r(x_{u},e_{d}))$ in Eq.(3) according to the BT model used in RLHF and substituting term $r(x_{u},e)$ in Eq.(3) with Eq.(6), the last two phases of RLHF pipeline can be equivalently transformed into optimizing DPO loss below:
$\mathcal{L}_{\rm DPO}=-\mathbb{E}_{(x_{u},e_{p},e_{d})}\left[{\rm log}\,\sigma\left(\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}\right)\right],$ (7)
wherein $\sigma(x)$ is the sigmoid function.
DPO is able to directly extract the optimal policy from pairwise preference data, making it more practical for preference alignment than RLHF. Nevertheless, DPO and RLHF are usually designed for pairwise preference. The oversight of other negative items impedes the performance of the LM-based recommenders. To bridge the gap, we expand DPO to S-DPO in recommendation tasks, in consideration of multiple negative items.
3 Methodology
3.1 Derivation of S-DPO loss
To align LM-based recommender $\mathcal{M}_{\theta}$ with multi-negative preference, we first derive the preference distribution and then propose a new loss function called S-DPO (depicted in Figure 1).
Multi-negative Preference Distribution.
As mentioned in Section 2, for user $u$, there is a partial ranking stipulating $i_{p}>_{u}i_{d},\forall\,i_{d}\in\mathcal{I}_{d}$ in sequential recommendation tasks. Let $\mathcal{E}_{d}$ be the titles of dispreferred items $\mathcal{I}_{d}$. The aforementioned partial ranking is equivalent to $e_{p}\succ e_{d}|x_{u},\forall e_{d}\in\mathcal{E}_{d}$, from which a multi-negative preference dataset $\{x_{u},e_{p},\mathcal{E}_{d}\}$ can be curated in an analogous way to RLHF.
For the dataset pairing one preferred item with multiple dispreferred items, we leverage the Plackett-Luce (PL) model [40, 41] to build preference distribution. Given prompt $x_{u}$, $K$ titles $e_{1},e_{2},\cdots,e_{K}$ and a permutation $\tau:[K]\rightarrow[K]$ reflecting user preference, with $\tau(j)$ denoting the $j$-th element of permutation $\tau$, the PL model estimates that the ranking $e_{\tau(1)},e_{\tau(2)},\cdots,e_{\tau(K)}$ turns out true, as:
$p(\tau|e_{1},e_{2},\cdots,e_{K},x_{u})=\prod_{j=1}^{K}\frac{{\rm exp}\left(r(x_{u},e_{\tau(j)})\right)}{\Sigma_{l=j}^{K}{\rm exp}(r\left(x_{u},e_{\tau(l)})\right)}.$ (8)
By enumerating all the permutations starting with $p$ and calculating sum of their probability given by the PL model, the final multi-negative preference distribution $p^{*}$ can be derived as:
$p^{*}(e_{p}\succ e_{d},\forall e_{d}\in\mathcal{E}_{d}|x_{u})=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}.$ (9)
For brevity, the complete derivation is delegated to Appendix A.1.
Deriving S-DPO.
By substituting reward function $r(x_{u},e)$ in Eq.(9) with Eq.(6), the mutli-negative preference distribution can be rewritten as:
$p^{*}(e_{p}\succ e_{d},\forall\,e_{d}\in\mathcal{E}_{d}|x_{u})=\frac{1}{1+\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}\left(\beta\,{\rm log}\,\frac{\pi(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}-\beta\,{\rm log}\,\frac{\pi(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}\right)}.$ (10)
Through plugging distribution given by Eq.(10) in the reward learning objective in Eq.(3), our S-DPO loss can be formulated for policy $\pi_{\theta}$ as:
$\mathcal{L}_{\rm S-DPO}(\pi_{\theta};\pi_{\rm ref})=-\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})\sim\mathcal{D}}\left[{\rm log}\,\sigma\left(-{\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}\left(\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}\right)\right)\right].$ (11)
Notably, when the number of candidates $N$ is 2, which means there is only one dispreferred item, S-DPO reduces to DPO. The proof is provided in Appendix A.2.
Gradient Analysis.
We conduct gradient analysis on S-DPO. The gradient of $\mathcal{L}_{\rm S-DPO}$ with respect to parameters $\theta$ takes the following formulation:
$\nabla_{\theta}\mathcal{L}_{\rm S-DPO}(\pi_{\theta};\pi_{\rm ref})=$
$-\beta\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\bigg{[}\underbrace{\sigma\left({\rm log}\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e_{d},e_{p},x_{u}))\right)}_{\text{higher weight when reward deviates from preference}}\cdot\bigg{[}\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{p}|x_{u})-\sum_{e_{d}\in\mathcal{E}_{d}}\frac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{d}|x_{u})}{\underbrace{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}_{\text{higher weight when reward is larger}}}\bigg{]}\bigg{]},$
wherein $g(e_{j},e_{k},x_{u})=r_{\theta}(x_{u},e_{j})-r_{\theta}(x_{u},e_{k})$ and similar to DPO, $r_{\theta}(x_{u},e)=\beta{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ is the implicit reward function defined by $\pi_{\theta}$. See Appendix A.3 for a complete derivation.
Recap the DPO gradient below:
$\nabla_{\theta}\mathcal{L}_{\rm DPO}(\pi_{\theta};\pi_{\rm ref})=-\beta\mathbb{E}_{(x_{u},e_{p},e_{d})}\bigg{[}\underbrace{\sigma(g(e_{d},e_{p},x_{u}))}_{\text{higher weight when reward is wrong}}\cdot\left[\nabla_{\theta}\,{\rm log}\pi_{\theta}(e_{p}|x_{u})-\nabla_{\theta}\,{\rm log}\pi_{\theta}(e_{d}|x_{u})\right]\bigg{]}.$
Similar to DPO, the gradient of S-DPO loss increases the likelihood of the preferred item and decreases the likelihood of all the dispreferred items. Each example is also weighed by how much the implicit reward $r(x_{u},e)$ deviates from the preference data. However, compared with DPO, S-DPO harnesses information of multiple dispreferred items in this weight.
Moreover, S-DPO contains an extra weight term only assigned to the gradient of dispreferred items. The term $\frac{1}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}=\frac{{\rm exp}(r_{\theta}(x_{u},e_{d}))}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(r_{\theta}(x_{u},e^{\prime}_{d}))}$ mirrors the relative reward of item $i_{d}$ in dispreferred items. The larger the reward of item $i_{d}$ is compared with other dispreferred items, the higher the weight will be. Therefore, for dispreferred items with larger reward, which can be considered as hard negative items, their likelihood will decline more in the next update. This mechanism endows S-DPO with more effectiveness and stability than DPO. The gradient of S-DPO also tends to be larger, enabling S-DPO to converge faster .
3.2 Properties of S-DPO
In this section, we will discuss the structural correlation between DPO and BPR [42], together with S-DPO and softmax loss [38], which demonstrates the advantage of S-DPO over DPO and language modeling objective.
For user $u$, preferred item $i_{p}$ and one dispreferred $i_{d}\in\mathcal{I}_{d}$, BPR loss takes the form:
$\mathcal{L}_{\rm BPR}=-\mathbb{E}_{(u,i_{p},i_{d})}\left[{\rm log}\,\sigma\left(f(u,i_{p})-f(u,i_{d})\right)\right],$ (12)
wherein $f(u,i)$ represents preference score of user $u$ for item $i$.
Similarly, given dispreferred item set $\mathcal{I}_{d}$, the softmax loss takes the form:
$\mathcal{L}_{\rm softmax}=-\mathbb{E}_{(u,i_{p},\mathcal{I}_{d})}\left[{\rm log}\,\sigma\left(-{\rm log}\sum_{i_{d}\in\mathcal{I}_{d}}{\rm exp}\left(f(u,i_{d})-f(u,i_{p})\right)\right)\right].$ (13)
Review the DPO loss in Eq.(7) and S-DPO loss in Eq.(11). Notably, term $\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ is the implicit reward function, denoted by $r_{\theta}(x_{u},e)$ in Section 3.1. According to Section 2, $r_{\theta}(e,x_{u})$ reflects the preference of user $u$ to item $i$ corresponding to title $e$. When the reference model has no knowledge about recommendation, i.e., $\pi_{\rm ref}(e|x_{u})$ is approximately a uniform distribution, term $r_{\theta}(x_{u},e)=\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ exactly reveals absolute preference. Hence, $r_{\theta}(x_{u},e)$ possesses a similar function to $f(u,i)$
From this perspective, DPO and S-DPO can be seen as special patterns of BPR and softmax loss, respectively. The structural correlation indicates that DPO and S-DPO are more suitable for recommendation than the language-modeling loss. Moreover, as softmax loss works better than BPR loss in multi-negative scenarios [19], it can be inferred that S-DPO will be more tailored for multi-negative user preference alignment than DPO.
4 Experiments
In this section, we aim to answer the following research questions:
  • RQ1: How does S-DPO compare with traditional and LM-based sequential recommendation models on performance?
  • RQ2: How does the LM-based recommender benefit from the multiple negatives?
Baselines.
We thoroughly compare S-DPO with three categories of recommenders in sequential recommendations: traditional recommenders (GRU4Rec [45], Caser [46], SASRec [47]), LM-enhanced recommenders (MoRec [48]) and LM-based recommenders (LLaMA2 [31], Chat-REC [14], TALLRec [13], LLaRA [18]). See detailed introduction and comparison of baselines in Appendix B.
Datasets.
We conduct extensive experiments on three real-world benchmark datasets which differ in size and domain (Movielens [49], Goodreads [^1] https://www.goodreads.com, and LastFM [50]). Following standard settings of [14, 18], we employ a commonly used metric Hit Ratio@1 (HR@1) for performance evaluation and an additional metric Valid Ratio to evaluate the LM-based methods’ ability to generate appropriate responses. See detailed introductions of datasets and evaluation metrics in Appendix B.
Implementation.
We implement all LM-based recommenders on 4 NVIDIA A100 GPUs. For all LM-based recommenders, we conduct a supervised fine-tuning stage for a maximum of 5 epochs. For S-DPO and its variants, we conduct a preference alignment stage for further 3 epochs. Different from existing methods, we only optimize loss on item titles and find it effective in recommendation tasks. Refer to Appendix B for more implementation details.
4.1 Overall Performance Comparison (RQ1)
Table 1 presents a comparative analysis of the performance of our proposed S-DPO and baselines. “Rel.Ipv” denotes the relative improvement of S-DPO compared with baselines. Bold and underlined indicate the best and the second-best performance, respectively. We observe that:
  • LM-based recommenders have driven impressive performance breakthroughs compared with traditional recommenders. Our results reveal that traditional recommenders outperform untuned LM-based recommenders (LLaMA, ChatRec) but fall short compared to LM-based recommenders fine-tuned on historical interactions (TALLRec and LLaRA). It is noted that untuned LM-based recommenders are limited by inadequate instruction-following capabilities or a lack of domain-specific knowledge indicated by the low valid ratio and suboptimal performance, which highlights the necessity of the supervised fine-tuning stage to further ground the inherent ability of language models down to sequential recommendation tasks. Moreover, MoRec also exhibits suboptimal performance compared to its traditional variant because it leaves the reasoning ability of LM untouched. The superior performance of recent LM-based recommenders indicates the significant roles of knowledge and reasoning ability in language models for recommendation tasks, which highlights the potential of LM-based recommenders.
  • Tailoring language model for recommendation task further boosts the performance of LM-based recommenders. For LM-based recommenders, the substantial performance gap between fine-tuned and untuned approaches emphasizes the importance of tailoring models for recommendations. TALLRec adapts LM for recommendation by supervised fine-tuning LM on historical interactions, surpassing traditional recommenders. Additionally, LLaRA consistently outperformed TALLRec across all datasets, suggesting that introducing collaborative signals through appropriate item representations is a viable direction for further adapting LM. However, existing LM-based methods adapt LM from either item representation methods or corpus construction, leaving the adaptation of optimization objectives unexplored. Instead, S-DPO aligns the language model with multi-negative user preference data by extending DPO to include a softmax ranking loss, making it a more appropriate loss function for recommendation tasks.
  • S-DPO consistently outperforms all traditional recommenders and the state-of-the-art LM-based recommenders on all datasets. S-DPO shows an improvement ranging from 11.10% to 47.03% on Hit Ratio@1 compared to the second-best baseline. Building on a supervised fine-tuning stage, we attribute this further improvement to the preference alignment stage, which explicitly instills ranking information into LM and utilizes preference data with multiple negative samples. Such superior performance suggests that explicitly tailoring LM for recommendation using user preference data at the training objective level is more effective than other LM-based recommenders. By leveraging the inherent abilities of the LM and incorporating ranking information from user preference data, S-DPO effectively differentiates between preferred and less preferred items. Notably, the preference alignment stage hardly harms the inherent ability of LM, illustrated by a high valid ratio.
Table 1: The performance comparison on three real-world datasets. The improvement achieved by S-DPO is significant ($p$-value <<much-less-than$<$ < 0.05).
LastFM Goodreads MovieLens
HR@1 ValidRatio Rel.Ipv HR@1 ValidRatio Rel.Ipv HR@1 ValidRatio Rel.Ipv
Traditional GRU4Rec [45] 0.3867 1.0000 70.91% 0.2616 1.0000 153.36% 0.3750 1.0000 40.35%
Caser [46] 0.4174 1.0000 58.34% 0.2233 1.0000 196.82% 0.3861 1.0000 36.31%
SASRec [47] 0.3581 1.0000 84.56% 0.2233 1.0000 196.82% 0.3444 1.0000 52.82%
LM-based LLaMA2 [31] 0.0233 0.3845 2736.48% 0.0246 0.3443 2594.31% 0.0421 0.4421 1150.12%
ChatRec [51] 0.3306 1.0000 99.91% 0.3770 1.0000 75.81% 0.2000 0.9895 163.15%
MoRec [48] 0.2877 1.0000 129.72% 0.1652 1.0000 301.21% 0.2822 1.0000 86.50%
TALLRec [13] 0.4983 0.9573 32.63% 0.4180 0.9836 58.56% 0.3895 0.9263 35.12%
LLaRA [18] 0.5292 0.9950 24.89% 0.4508 0.9918 47.03% 0.4737 0.9684 11.10%
Ours S-DPO 0.6609 0.9900 - 0.6628 0.9992 - 0.5263 0.9895 -
4.2 Study on S-DPO (RQ2)
Ablation Study.
(a) Ablation study.
(b) Study of validation loss.
(c) Study of preferred item reward.
Figure 2: Study on S-DPO. (2(a)) Ablation study of S-DPO compared with SFT and DPO on three datasets. (2(b)) Comparison of the trend of validation loss between DPO and S-DPO on LastFM. (2(c)) Comparison of the reward of preferred items between DPO and S-DPO on LastFM.
To investigate the effect of explicit ranking optimization and multiple negative samples of S-DPO, we compare it with Supervised Fine-Tuning(SFT), and a variant of S-DPO with only a single negative sample, downgrading to pairwise DPO loss. The experimental results are reported in Figure 2(a). We can observe that DPO can achieve an overall better performance compared to SFT, which underscores the effectiveness of instilling ranking relationships into existing LM-based recommenders. With a more effective ranking gradient provided by multiple negative samples, S-DPO can further boost performance and achieve the best among all baseline methods and variants.
(a) Study of negative samples.
(b) Study of $\beta$ on Hit Ratio@1.
(c) Study of $\beta$ on Valid Ratio.
Figure 3: Studies on values of $\beta$ and negative samples numbers of S-DPO on LastFM. (3(a)) Performance comparisons with varying numbers of negative samples ($\beta$ = 1). (3(b)) Performance comparisons with varying values of $\beta$ setting negative samples number as 3. (3(c)) Validity comparisons with varying values of $\beta$ setting negative samples number as 3.
Study on number of negative samples.
Benefiting from the utilization of multiple negative pairs in preference data, our S-DPO offers two empirically appealing properties compared to DPO: 1) S-DPO has more effective gradients facilitating the optimization; 2) S-DPO mitigates the data likelihood decline issue of DPO. Figure 2(b) provides the comparison of validation loss between S-DPO and DPO, illustrating that the loss of S-DPO decreases faster and more significantly. This observation demonstrates that multiple negative pairs provide larger and more meaningful gradients for model optimization, which is attributed to the inherent ability of S-DPO to mine negative samples [37] (cf. Section 3.1). More results of loss analysis can be found in Appendix C.
On the other hand, it is widely accepted that the optimization of DPO suffers from data likelihood decline issues [52, 53], which implies that the log-likelihood of the preferred completions is reduced below the original log-likelihood from the reference model, hindering the performance of DPO. We, therefore, study the behavior of S-DPO and surprisingly find that it has the potential property of mitigating data likelihood decline issues. As illustrated in Figure 2(c), S-DPO exhibits continually increasing rewards of preferred items, while DPO struggles to increase the reward of preferred items.
To further verify the superiority of the multiple negative samples of S-DPO compared with DPO, we conduct experiments to explore the potential of the number of negative samples, with the results depicted in Figure 3(a). It can be observed that utilizing multiple negative samples allows the model to achieve better performance than with a single one. Furthermore, as the number of negative samples increases, the model’s performance exhibits continual improvements. We attribute this success of S-DPO to more effective ranking gradients provided by multiple negatives which can be connected to the superior performance of contrastive loss in self-supervised recommendations [37, 38, 54]. More results of preferred item rewards can be found in Appendix C.
Study on values of $\beta$.
In S-DPO, $\beta$ is a hyperparameter controlling the deviation of LM from the base reference policy [33]. Typically, a smaller value of $\beta$ implies that the language model is more heavily influenced by the preference signals and vice versa. In this section, we explore the effect of $\beta$ on S-DPO. As indicated in Figure 3(b) through 3(c), a higher $\beta$ can achieve overall better performance in our task, while a lower $\beta$ may overwhelm the model’s learned knowledge from the supervised fine-tuning stage, as evidenced by both low valid ratio and hit ratio. On the other hand, an excessively large $\beta$ prevents the model from effectively learning ranking relationships, leading to suboptimal performance. In all our main experiments and studies, we set $\beta$ as 1 to achieve a balance between ranking signals and inherent knowledge of language models.
5 Related Work
5.1 LM for Recommendation
Recent advancements in recommendation systems have increasingly incorporated Language Models (LMs) due to their extensive knowledge and robust reasoning abilities. This integration occurs primarily in two forms: LM-enhanced recommenders and LM-based recommenders. LM-enhanced recommenders utilize LM embedding as semantic representations to provide contrastive signals [55, 56, 57] or utilize LM as advanced feature extractors improving the representation of user and item features [58, 59, 60]. However, these systems still rely on traditional recommenders for the final recommendation task, which leaves the reasoning ability of LM largely untouched.
On the other hand, LM-based recommenders directly employ LMs for making recommendations. Early works leverage LMs’ in-context learning capabilities for zero-shot or few-shot recommendations, demonstrating significant potential [9, 10, 11, 14]. However, untuned LM-based recommenders are limited by inadequate instruction-following capabilities and a lack of domain-specific knowledge. To bridge this gap, recent efforts in this category include supervised fine-tuning of LMs on the historical interactions to enhance their performance in recommendation tasks [13, 16, 22, 23, 26]. More recently, researchers have discovered that exploring item representation methods in the finetuning phase may further boost LM’s ability for recommendation [29]. This branch of works includes integrating collaborative signals [17, 18, 24, 61], adjusting numeric representations [21, 62, 63] or introducing additional item tokens [15, 25, 27, 28].
However, existing finetuned methods follow the training objective of language generation without any specific adjustments for personalized ranking. Different from them, S-DPO proposes to explicitly optimize item ranking information on preference data.
5.2 Preference Alignment of Language Models
Reinforcement Learning from Human Feedback (RLHF) [30, 31, 32] is a prevalent method of LMs to learn from human preferences. The RLHF pipeline comprises reward model learning and reinforcement learning (RL) optimization, the latter of which suffers instability and inefficiency. Direct Preference Optimization (DPO) [33] bypasses the brittle RL phase via a particular reward model parameterization and is thus simpler to implement while still keeping the performance of RLHF.
DPO proves to be effective in many scopes, like NLP [33, 64] and multimodal LMs [34, 65, 66, 67]. Besides, several variants have been proposed for further improvement of DPO. $\Psi$PO [68] is a generalization of DPO loss and its representative IPO can better overcome the problem of overfitting. ODPO [35] treats preference pairs differently by stipulating that the likelihood gap of two responses should be greater than a corresponding offset value. Other variants like KTO [69], $f$-DPO [70], RSO [71], etc. also enhance DPO in various aspects. Despite these contributions, the possibilities for leveraging and further adapting DPO for recommendation are still largely unexplored. Moreover, there are few studies that discuss extending DPO to handle multi-negative scenarios.
6 Limitation
Despite effectiveness, there are several limitations not addressed in this paper. On the one hand, the number of negative samples is capped at 15 in our experiments. The potential of multiple negative samples hasn’t been fully explored due to the limited time and computation resources. On the other hand, increasing the number of negative examples inevitably results in higher training costs, a phenomenon that becomes more pronounced as the number of negative examples grows in the context of language models. Finally, despite the empirical success and theoretical connection to softmax loss in recommendation tasks, the profound understanding of softmax ranking loss on LM needs to be further explored.
7 Conclusion
In this work, we devised a principled Softmax-DPO (S-DPO) loss specially tailored for LM-based recommenders, utilizing multiple negatives in preference data to explicitly instill ranking information into LM. Empirically, S-DPO surpasses all baseline models including traditional and LM-based methods on three datasets in sequential recommendation tasks while successfully mitigating the data likelihood decline issue of DPO. Grounded by theoretical proof, we bridge S-DPO with the softmax loss in self-supervised recommendations, underscoring the ranking performance of S-DPO and highlighting the critical roles of multiple negatives. Also, we theoretically find that S-DPO has an inherent ability to mine hard negatives which provide larger and more effective gradients to model optimization, assuring its exceptional capabilities in recommendation tasks. We believe that our S-DPO, as a generalization of DPO, provides valuable insights for future LM-based recommenders and has the potential to benefit research fields other than recommender systems [^2] Thc extension to broader impact will be detailed in Appendix D.
Appendix A Mathematical Derivations
A.1 Deriving Preference Distribution
The PL model takes the form:
$p^{*}(\tau|e_{e},i_{2},\cdots,e_{K},x_{u})=\prod_{j=1}^{K}\frac{{\rm exp}\left(r(x_{u},e_{\tau(j)})\right)}{\Sigma_{l=j}^{K}{\rm exp}(r\left(x_{u},e_{\tau(l)})\right)}.$ (14)
The ranking in multi-negative preference data is $e_{p}\succ e_{d}|x_{u},\forall e_{d}\in\mathcal{E}_{d}$. Our new preference distribution that estimates the probability of the ranking can be derived:
$p^{*}(e_{p}\succ e_{d},\forall\,e_{d}\in\mathcal{I}_{d}|x_{u},e_{p},\mathcal{E}_{d})$ $=\sum_{\tau\in\{\tau^{\prime}|\tau^{\prime}(1)=p\}}p^{*}(\tau|x_{u},e_{p},\mathcal{E}_{d})$
$=\sum_{\tau\in\{\tau^{\prime}|\tau^{\prime}(1)=p\}}\prod_{j=1}^{K}\frac{{\rm exp}\left(r(x_{u},e_{\tau(j)})\right)}{\sum_{l=j}^{K}{\rm exp}\left(r(x_{u},e_{\tau(l)})\right)}$
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in Per(ind(\mathcal{I}_{d}))}\prod_{j=1}^{K-1}\frac{{\rm exp}\left(r(x_{u},e_{\tau^{\prime}(j)})\right)}{\sum_{l=j}^{K-1}{\rm exp}(r(x_{u},e_{\tau^{\prime}(l)}))}$
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}(r(x_{u},e_{j}))}\times\sum_{\tau^{\prime}\in Per(ind(\mathcal{I}_{d}))}p^{*}(\tau^{\prime}|x_{u},\mathcal{E}_{d})$
$=\frac{{\rm exp}(r(x_{u},e_{p}))}{\sum_{j=1}^{K}{\rm exp}\left(r(x_{u},e_{j})\right)},$
wherein $ind(\mathcal{E}_{d})$ denotes the indices of titles in $\mathcal{E}_{d}$ and $Per(ind(\mathcal{E}_{d}))$ denotes the set of permutations of index set $ind(\mathcal{E}_{d})$. The third equation is because a permutation of $\{1,2\cdots,K\}$ starting with $p$ can be divided into the prefix $p$ and a subsequent permutation of the rest indices, which is exactly $Per(ind(\mathcal{E}_{d}))$.
A.2 Connection Between DPO and S-DPO
When $N=2$, the following equations hold:
$\mathcal{L}_{\rm S-DPO}(\pi_{\theta};\pi_{\rm ref})$ (Eq.(11))
$=-\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[{\rm log}\,\sigma\left(-{\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}\left(\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}\right)\right)\right]$
$=-\mathbb{E}_{(x_{u},e_{p},e_{d})}\left[{\rm log}\,\sigma\left(-{\rm log}\,{\rm exp}\left(\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}\right)\right)\right]$ ($N=2$)
$=-\mathbb{E}_{(x_{u},e_{p},e_{d})}\left[{\rm log}\,\sigma\left(\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}\right)\right]$
$=\mathcal{L}_{\rm DPO}(\pi_{\theta};\pi_{\rm ref}).$
Therefore, DPO is a special case of S-DPO.
A.3 Deriving the Gradient of S-DPO Loss
Let $V(\theta;e_{d})=g(e_{d},e_{p},x_{u})=\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}-\beta\,{\rm log}\,\frac{\pi_{\theta}(e_{p}|x_{u})}{\pi_{\rm ref}(e_{p}|x_{u})}$ and the S-DPO loss takes the following form:
$\mathcal{L}_{\rm S-DPO}(\pi_{\theta};\pi_{\rm ref})$
$=-\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[{\rm log}\,\sigma\left(-{\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right)\right]$ (15)
The gradient of $V(\theta;e_{d})$ can be formulated as:
$\nabla_{\theta}\,V(\theta;e_{d})=\beta(\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{d}|x_{u})-\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{p}|x_{u}))$ (16)
Using properties of sigmoid function that $\sigma^{\prime}(x)=\sigma(x)(1-\sigma(x))=\sigma(x)\sigma(-x)$ and thus $\left(({\rm log}\,\sigma(x)\right)^{\prime}=\frac{1}{\sigma(x)}\times\sigma(x)\sigma(-x)=\sigma(-x)$, we have:
$\nabla_{\theta}\,\mathcal{L}_{\rm S-DPO}(\pi_{\theta};\pi_{\rm ref})$
$=-\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\nabla_{\theta}\,{\rm log}\,\sigma\left(-{\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right)\right]$
$=\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\sigma\left({\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right)\cdot\nabla_{\theta}\,{\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right]$ ($\left({\rm log}\sigma(x)\right)^{\prime}=\sigma(-x)$)
$=\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\sigma\left({\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right)\cdot\frac{\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\cdot\nabla_{\theta}\,V(\theta;e_{d})}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e^{\prime}_{d}))}\right]$
$=-\beta\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\sigma\left({\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))\right)\cdot\sum_{e_{d}\in\mathcal{E}_{d}}\frac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{p}|x_{u})-\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{d}|x_{u})}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e^{\prime}_{d})-V(\theta;e_{d}))}\right]$ (Eq. (16))
$=-\beta\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\sigma\left({\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e_{d},e_{p},x_{u}))\right)\cdot\sum_{e_{d}\in\mathcal{E}_{d}}\frac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{p}|x_{u})-\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{d}|x_{u})}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}\right]$ (Definition of $V(\theta;e_{d})$)
$=-\beta\mathbb{E}_{(x_{u},e_{p},\mathcal{E}_{d})}\left[\sigma\left({\rm log}\,\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e_{d},e_{p},x_{u}))\right)\cdot\left[\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{p}|x_{u})-\sum_{e_{d}\in\mathcal{E}_{d}}\frac{\nabla_{\theta}\,{\rm log}\,\pi_{\theta}(e_{d}|x_{u})}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}\right]\right]$
The last equation is because:
$\sum_{e_{d}\in\mathcal{E}_{d}}\frac{1}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}=\frac{\sum_{e_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e_{d}))}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(V(\theta;e^{\prime}_{d}))}=1$
Appendix B Experimental Settings
B.1 Baselines
We compare the performance of S-DPO, against both traditional and LM-based baselines to showcase the effectiveness of our method. Specifically, for traditional methods, we have:
  • GRU4Rec [45] utilizes the GRU (Gated Recurrent Unit) architecture to model sequences, enabling effective prediction in recommendation tasks.
  • Caser [46] employs both horizontal and vertical convolutional operations to enhance the capture of high-order interactions within item sequences, improving recommendation accuracy.
  • SASRec [47] incorporates a multi-head self-attention mechanism in its self-attentive sequential recommendation model, facilitating the modeling of intricate sequential data patterns.
For LM-enhanced method, we have:
  • MoRec [48] advances traditional recommendation systems by incorporating the modality features of items instead of the id feature. we employ BERT for the text encoder and SASRec for the recommendation architecture.
For LM-based methods, we have:
  • LLaMA2 [31] utilized vanilla LLaMA2-7B to directly generate recommendation results through direct prompting.
  • Chat-REC [14] is implemented based on the framework discussed in [14], we retain user interaction sequences consisting of item titles as use profiles for a fair comparison. We use GPT4 [51] as its primary large language model.
  • TallRec [13] first propose to transform interaction sequences into textual prompts and then fine-tunes large language models using domain-specific corpus.
  • LLaRA [18] combines collaborative signals from traditional recommendation systems into the fine-tuning of large language models for improved recommendation performance.
B.2 Datasets
Table 2: Statistics of datasets.
Dataset MovieLens Goodreads LastFM
#Sequence 943 6,031 1,220
#Items 1,682 4,500 4,606
#Interactions 100,000 220,100 73,510
To evaluate the effectiveness of S-DPO, we conduct experiments on three widely used real-world datasets: Movielens [49], Goodreads [^3] https://www.goodreads.com, and LastFM [50]. The statistics of datasets are illustrated in Table 2. The MovieLens dataset is widely used for movie recommendation tasks and includes user ratings and movie titles, we select the MovieLens100K dataset in our experiment. Similarly, Goodreads is sourced from a social book cataloging website, where users can explore, rate, and review a variety of books. LastFM dataset comprises users’ listening history and artists’ names from the Last.fm online music platform. Following [18], we maintain their titles as textual descriptions for each dataset. For Goodreads, we remove users and books with less than 20 interactions, which keeps the same as the processing of MovieLens. For all datasets, we organize sequences chronologically before dividing the data into training, validation, and testing sets in an 8:1:1 ratio to prevent any potential information leakage.
B.3 Implementation Details
We implement all approaches with Python 3.9.7, PyTorch 2.2.2, and transformers 4.39.0 on 4 NVIDIA A100s. We select Llama2-7B [31] as the LM backbone for S-DPO. Following [18], we randomly select prompts from several prompt formats during training and evaluation to ensure flexibility and generality. For optimization of all the traditional methods, the Adam optimizer is employed with a learning rate adjusted to 0.001, and a batch size configured at 256. All models undergo L2 regularization, with coefficients experimentally determined from [1e-3, 1e-4, 1e-5, 1e-6, 1e-7]. In all experiments involving large language models, we train each method for a maximum of 5 epochs using a batch size of 128 and select the checkpoint with the lowest loss on the validation set as the final checkpoint. A warm-up strategy is applied to the learning rate, starting at 5% of its maximum value, and gradually adjusting it through a cosine scheduler throughout the training process. For S-DPO and all of its ablation studies, we further conduct preference training for further 3 epochs with a batch size of 128 and a learning rate of 1e-5. Setting the value of $\beta$ as 1, we search the number of negative samples in [3,5] for the main results. The effects of both factors are further explored in 4.2.
B.4 Evaluation Metrics
Given that LMs primarily produce textual responses rather than comprehensive item rankings, we utilize a re-ranking metric in line with previous research [18] to assess recommendation performance. For each sequence, a candidate set is constructed by randomly selecting 20 non-interacted items and always includes the correct item. We assess all models based on their ability to pinpoint the correct item within this candidate set, employing the HitRatio@1 (HR@1) metric for performance evaluation. Following [18], we also introduce an additional metric called the Valid Ratio to evaluate the LM-based methods’ adherence to instructions and their ability to generate appropriate responses. Due to the difficulty LMs face in producing ranked results for candidate items, position-aware metrics like NDCG are deemed unsuitable for this evaluation.
Appendix C Study of Negative Samples Number
(a) Study of valid loss on MovieLens.
(b) Study of valid loss on Goodreads.
Figure 4: Study on validation loss between S-DPO and DPO on MovieLens and Goodreads.
Figure 4(a) through figure 4(b) provide the comparison of validation loss between S-DPO and DPO on MovieLens and Goodreads, illustrating that the loss of S-DPO decreases faster and more significantly. This observation is consistent with Figure 2(b).
On the other hand, figure 5(a) through figure 5(b) showcase that S-DPO exhibits continually increasing rewards of preferred items, while DPO struggles to increase the reward of preferred items on both MovieLens and Goodreads, which is aligned with Figure 2(c).
(a) Study of preferred item rewards on MovieLens.
(b) Study of preferred item rewards on Goodreads.
Figure 5: Study on validation loss between S-DPO and DPO on MovieLens and Goodreads.
Appendix D Broader Impact
We left further exploration of softmax ranking loss in LM including more negative samples and validation on various settings as future works. We believe that S-DPO, a generalization of DPO loss has the potential to benefit other research areas other than recommender systems. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.